OpenStudy (anonymous):

Simplify: radical 18 divided by radical 15

OpenStudy (anonymous):

\[\frac{ \sqrt{18} }{ \sqrt{15}}\]

OpenStudy (anonymous):

Yep that's the problem. :)

OpenStudy (anonymous):

You have to factor out the numbers! So factor 18 first...

OpenStudy (anonymous):

Well...I don't really know how

OpenStudy (anonymous):

factoring is basically listing the numbers that you multiply to get the main number..

OpenStudy (anonymous):

so what two numbers multiply to get 18?

OpenStudy (anonymous):

2 and 9

OpenStudy (anonymous):

mhm...

OpenStudy (anonymous):

So what two numbers multiply to get 9?

OpenStudy (anonymous):

3 and 3

OpenStudy (anonymous):

so 18 factored is 2*3*3

OpenStudy (anonymous):

so when simplfying radicals you find the numbers that are there twice. What number is in the list twice?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

okay so you take that 3 and put it outside the radical so it becomes...\[3\sqrt{2}\]

OpenStudy (anonymous):

K well what about 15 because the numbers don't happen twice

OpenStudy (anonymous):

Right so its simplified all the way already.

OpenStudy (anonymous):

Oh...Okay so now what?

OpenStudy (anonymous):

Its simplified!

OpenStudy (anonymous):

https://www.khanacademy.org/math/arithmetic/exponents-radicals/radical-radicals/v/simplifying-radicals check this out for a detailed verson

OpenStudy (anonymous):

Oh um well my teacher gives us the answers and he put radical 30 divided by 5.

OpenStudy (anonymous):

???